When an elastic material with Youngs modulus Y is subjected to stretching stress S, elastic energy stored per unit volume of the material is  


        (a)    YS/2


        (b)    S2Y/2


        (c)    S2/2Y        


        (d)    S/2Y


Grade:10

1 Answers

Amit Saxena
10 years ago

(c)
Energy stored per unit volume
    =  1/2×  stress×  strain
    =  1/2×  stress× (stress/Young''s modulus)    
    =1/2×(stress)^2/((Young^'' s modulus) )=S^2/(2 Y)